OpenStudy (anonymous):

Find the integral of dt/t from x to 2x.

OpenStudy (anonymous):

\[\int_x^{2x}\frac{dt}{t}\]?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

what is the anti deriviative of \(\frac{1}{t}\) ?

OpenStudy (anonymous):

That's the part I got stucked.

OpenStudy (anonymous):

ok it is the log

OpenStudy (anonymous):

How do I find that?

OpenStudy (anonymous):

you don't "find" it, you know it \[\frac{d}{dt}[\ln(t)]=\frac{1}{t}\]

OpenStudy (anonymous):

So it's ln2x-lnx, right? And then it's ln(2x/x)=ln2, right?

OpenStudy (anonymous):

yeah i guess so

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

interesting that it is a constant, right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

\[\int_5^{10}\frac{dt}{t}=\ln(2)\]

OpenStudy (anonymous):

How about the integral of cos t dt from 1 to x? Isn't it sin(x)-sin(1)?

OpenStudy (anonymous):

@satellite73